Welcome >  Stations >  Greenbelt, Maryland (7105) >  Station Logs

Station History Log - Greenbelt, Maryland (7105)

SOD Year Day
of Year
Time
of Day
Data Impact System Description
71050725 2022 239 06:50 1 6.01.01 The ARSU (G77944) was replaced with ARSU (001204) on 08/27/2022 at 10:00am. The system changed from 85.92ns to 86.94ns
71050725 2022 236 06:34 1 6.01.01 The system was optimize, laser setting with the replacement of the power units PU610 and PU620 and ran into a few issues: The delay adjustments are now made on the individual capacitor banks. The CB631 was adjusted to 300usec (as before), and the CB632 delay was adjusted obtain maximum amplifier IR power. The delay for the CB632 is 190 usec. Adjusted oscillator PU610 voltage 50 volts above oscillator lasing threshold. The new voltage is 1.24kv. I noticed the oscillator pulse train was unstable. Adjusted rear cavity mirror, AML brag angle and cavity length for optimal pulse width and stability. The PU620 amplifier voltage was adjusted to 1.42kv (seems high) to obtain a 2.0 watts reading at 10PPS. Measured the green power at 0.74 watts (LOW!). Verified no self lasing, Pockel Cell fine adjustment and 1/4 waveplate adjustments were optimal. In stalled a replacement SHG to obtain 1.0 watts at 10pps. (50 % conversion) This is still a little low but is acceptable. Installing the replacement SHG shifted the angle of the beam at the output. Adjusted the final (21b) mirror to center the beam at the laser output. You will need to perform a laser bore site alignment. Ranged to the calibration target and took a data set. RMS was improved to 2.5 mm.
71050725 2022 183 10:00 1 6.01.01 The ARSU (001204) was replaced with ARSU (G77399) on 07/02/2022 at 10:00am. The system changed from 86.99ns to 85.94ns
71050725 2022 141 11:30 1 6.01.01 The ARSU (G77399) was replaced with ARSU (001204) on 05/21/2022 at 07:30am. The first tracking data received was on 05/23/2022 at 0443z, the system changed from 85.67ns to 86.98ns. The ARSU system delay changed from 150.80ns to 153.00ns
71050725 2022 117 02:50 1 6.01.01 The ARSU (001204) was replaced with ARSU (G77399) on 04/23/2022 at 07:00am. The first tracking data received was on 04/27/2022 at 0250z, the system changed from 86.93ns to 85.67ns
71050725 2022 020 14:00 1 6.01.01 Replaced the LRC board back to Moblas 7 original board, the one that was in the system was the network spare under test.
71050725 2021 356 14:00 1 6.01.01 The Moblas 7 ETM #001196 was replaced with the network spare ETM #001108 for testing. The system delay changed from 96.99 to 96.93
71050725 2021 320 13:00 0 99 The CRDv2 generation software was added to MOBLAS-7 for testing, in parallel with the operational CRD processes. The software was installed on Tuesday, 11/16/2021. There should be no change to operational processes at the station. The test data will be automatically sent from the station to the SLR-DOC, and will be examined by the DOC team for a month or so before moving into a prime mode.
71050725 2021 267 12:06 2 6.01.01 Corner Cube target "C" LTN 92-E was dropped in the grass and we could not locate it. The corner cube "B" LTN 92-F was installed on target "C" from 267 to 270. Corner cube LTN 94-1 was placed on target "C" on 270 and LTN 92-F was put back on target "B".
71050725 2021 223 04:55 0 12.01.02 the pressure comparisons. ("Met data additional information" - "Logging data from 2ndary Met unit 3x per day")
71050725 2020 288 11:11 0 99 The Dell new processor computer was installed into the system with CentOS 7
71050725 2020 231 19:20 2 6.01.01 It was determined that all cals and satellite voltage will be -2900 volts during the daylight hours..
71050725 2020 227 03:14 2 6.01.01 There was a meeting on 08/13/20 (226) to discuss Moblas 7 date accuracy, it was determined that all cals and satellite voltage will be -3100 volts. The system delay remained the same 87.06ns but the RMS is higher on the cals and some satellites. Before the voltage change cals 1.6mm after the voltage change now 4.1mm
71050725 2020 195 00:30 2 6.01.01 Replaced MCP PMT with the spare, do to weak returns. The system delay changed from 86.36ns to 87.04ns
71050725 2020 193 09:30 2 6.01.01 Replaced the high voltage cable, the old one had a bad connector. No change in the system delay.
71050725 2020 192 04:01 2 6.01.01 Replaced MCP PMT with the spare, do to weak returns. The system delay changed from 86.36ns to 87.04ns
71050725 2020 156 08:33 2 12.01.02 Replaced the MET 4 - Did not output humidity reading correctly
71050725 2020 146 02:22 0 6.01.01 The integrator cable W515 was replaced, from the Delay Box integrator output to the integrator input
71050725 2020 067 12:29 2 6.01.01 Changed the transmit cable with a low loss cable. The system changed from 83.38ns to 86.38ns.
71050725 2020 053 12:00 2 6.01.01 Test a PMT cable, signal reflections in the range window and to reduce high RMS.
71050725 2020 049 13:13 0 9.01.01 Replaced the GPS signal splitter with a better splitter for the GPS Antenna.
71050725 2020 043 11:25 2 99 Replaced the ARSU for testing. This is what changed the system delay by 1.25ns, from 82.05ns to 83.30ns.
71050725 2020 017 04:32 0 6.01.01 Changed HEO Etalons 0525 and 4146 also Glonass 9128-9139 from 4pps to 5pps
71050725 2019 243 03:45 1 9.02.03 ETN 2865 -Replacement of Xli with Stanford Research FS740 GPS steered Rubidium.
71050725 2019 208 12:00 2 6.01.01 Replaced PMT cable during troubleshooting to reduce high RMS. Awaiting analysis for a week or more.
71050725 2019 112 15:00 0 09:01:01 Removed G77979 CNS Clock and installed M04600 CNS Clock in the MOBLAS 7 Timing Rack.
71050725 2019 095 13:00 1 9.02.03 Removed XL-DC SCNS03102 and installed XLi SCNS000150 for testing.
71050725 2019 079 14:00 1 9.02.03 A second 1PPS output was connected to the LRC 1PPS. The prime 1PPS was connected to the ETM.
71050725 2019 077 18:00 2 12.01.02 MET 4 not updating - rebooted the controller all came back up operational.
71050725 2019 074 14:00 0 12.01.02 Swapped MET 4 SCNS001144 with SCNS000992.
71050725 2019 055 00:00 0 4.01 Replaced azimuth slip ring and three elevations slip rings due to mount cable failures.
71050725 2019 030 00:00 0 04.01 Elevation Tachometer had intermittent problems and was unable to perform calibrations.
71050725 2019 023 00:00 2 6.01.01 Reinstalled Assemblies Inc receive cable since the MegaPhase receive cable failed.
71050725 2018 324 00:00 2 9 A ND was added to the PD-10 start diode to reduce the signal on 11/20/18 (324) to try to eliminate the change from pre and post-cal system delay change, no help.
71050725 2018 282 04:50 2 06.01.03 Moblas 7 MCP was reinstalled max rating -3400
71050725 2018 265 12:45 2 06.01.03 A new MCP was installed max rating -3200
71050725 2018 250 16:30 2 99 modified LRC board installed for 10pps tracking of HEO
71050725 2018 240 00:00 0 99 Processor computer was upgraded with CentOS 6 vice RedHat operating system.
71050725 2018 198 17:30 2 6.01.01 Removed original Assemblies Inc. receive cable and replaced with Megapath receive cable.
71050725 2018 198 03:24 2 6 Installed new vendor (Megaphase) receive cable. Configuration flags changed, controller from 6 to 7 ad on the processor from 7 to 8.
71050725 2018 177 00:00 2 6.01.03 Found IRIG-B cable on ETM when it should have been connected to the ETC and XL-DC year was 8018 versus 2018 (reset to 2018).
71050725 2018 176 00:00 2 6.01.03 Bad ETM data reported - Tom V requested shut down of system.
71050725 2018 171 00:00 2 9 Replaced PD-10 with repaired unit - adjusted laser due to low power.
71050725 2018 169 00:00 2 9 Replaced failed PD-10 power supply for photo diode.
71050725 2018 164 00:00 2 9 PD-10 power supply for photo diode failed.
71050725 2018 108 00:00 1 6 System delay changed from 82.50ns to 82.55ns due to IRIS being partially closed.
71050725 2018 081 00:00 2 9 Replaced intermittent PD-10 power supply cable.
71050725 2018 078 00:00 2 9 Problems with intermittent PD-10 power supply cable - to be replaced at a later date.
71050725 2018 067 00:00 0 6.01.03 Tom O removed power dividers for the Event Timer receive NIM pulse to see if it would improve the ETM RMS - it did not. Tom V thinks it is a laser pulse-width problem.
71050725 2018 063 00:00 2 6.01.03 Tom O installed power dividers for the Event Timer receive NIM pulse onto a blank panel w/ground strap. Reconnected the fake stop and the HSLR NIM signal cables (extended by one nanosecond). Ground calibration RMS results for the TIU=3.6mm and the ETM=2.0mm.
71050725 2018 049 00:00 2 99 Removed the Moblas 8 ARSU that was being tested - installed the Moblas 7 ARSU.
71050725 2018 044 00:00 2 99 Installed an ARSU (for MOBLAS 8) for testing.
71050725 2018 043 00:00 2 9 ND`s were added to reduce the amplitude from .6 vns to .4 vns on the controller monitor. This improved the ETM RMS from 2.3mm to about 1.7mm.
71050725 2018 027 00:00 2 6 Gating cable stopped working - replaced on day 028.
71050725 2018 027 00:00 2 9 Installed new start diode mount and optimized for 10pps ranging.
71050725 2018 011 00:00 2 6 Intermittent gating cable, losing signal every so often.
71050725 2017 301 00:00 2 6 Replaced the received cable with a new one that is insulated, a stability test was performed. The system delay changed from 82.21ns to 82.62ns.
71050725 2017 299 00:00 2 99 Replaced the ARSU cable with a new cable.
71050725 2017 250 00:00 1 9:01:01 Remove the XL-DC under test and to reinstall MOBLAS 7 XLDC.
71050725 2017 240 13:00 1 9:01:01 Installed the M6 XL-DC for evaluation.
71050725 2017 216 00:00 0 99 Replaced the Processor computer with the one from the Lab.
71050725 2017 193 00:00 0 6 Reinstalled the MOBLAS 7 event timer 2394491/001196, the event timer 2394493/001108 was removed.
71050725 2017 163 00:00 2 9 NDs was added to the start diode to reduce the reading from 1.0vns to 0.5vns on the controller monitor to make the reading and laser power output comparable (100mjoules out = .5 vns. 06/12/17).
71050725 2017 140 00:00 2 6 Replaced the receive cable, the system delay changed from 84.66ns to 83.08ns.
71050725 2017 109 14:00 2 6.01.03 Event timer change from Cybi-ETM-02 s/n 1309-B-010 to s/n B-009
71050725 2017 109 14:00 0 6.01.03 Event timer change from Cybi-ETM-02 s/n 1309-B-010 to s/n B-009.
71050725 2017 099 00:00 0 99 Processor computer (DPC) was changed.
71050725 2017 087 00:00 1 12.01.02 MET4 was not inputting the data to the Controller (ICC), the power cable was disconnected in the rack, connected the cable and tie wraps were added to prevent this from reoccurring in the future.
71050725 2017 064 00:00 0 4 Completed the installation of the GLM Telescope, also balanced the MOBLAS 7 telescope.
71050725 2017 059 00:00 2 99 Changed the transmit delay from 2 to 1 on the LRC board.
71050725 2017 029 00:00 2 99 Installed repaired ARSU - caused system delays.
71050725 2016 252 00:00 1 6 Receive cable replaced.
71050725 2016 241 00:00 1 99 Spare LRC board installed into MOBLAS 7
71050725 2016 208 17:00 2 6 Event Timer went prime for operational support.
71050725 2016 149 00:00 2 12.01.02 Adjusted MET4 height so that the baraometer port is now level to the center of the elevation axis
71050725 2015 322 20:15 0 99 Engineering removed MOBLAS7 TXA card, ACD card, AZ encoder module and reference module and transferred them to MOBLAS-6 leaving MOBLAS-7 with severe AZ and EL glitching problems.
71050725 2015 321 13:32 1 6.01.02 Changed the HP5370B TIC from unit G78770 to unit G78769, the configuration flags was changed on the processor from 5 to 6.
71050725 2015 320 23:32 0 99 Changed correlation on AZ encoder from 4 to 3 and changed EL encoder from 4 to 2. Changed the AZ offset from +0.569 to -0.131.
71050725 2015 295 03:15 2 99 A new ephemeris file update for the star cal program was tested, the old ephemeris file will expire at the end of 2015.
71050725 2015 266 15:15 0 99 The Event Timer is now taking data at the same time as the 5370B counter and both set of data collection are processed at the same time.
71050725 2015 252 12:35 2 99 The 5370B counter that was under test was removed and shipped to Moblas 6, Moblas 7 5370B counter was reinstalled.
71050725 2015 247 10:50 2 99 The LRC board was exchanged with Moblas 7 LRC board for testing.
71050725 2015 215 11:15 0 99 Three cables for new servo system was added a 10Mhz and IRIG B from MOBLAS 7 XLDC and a 1pps from the LRC chassis, config flags were incremented (216 09:30) controller from 3 to 4 and processor from 4 to 5.
71050725 2015 215 10:40 2 99 Removed HP5370B counter G78681 and installed G78762 and the boards in the LRC chassis was changed to clear a transpwr reading on the controller monitor. Incremented the configuration flags (on day 216 03:25) on the controller from 3 to 4 and the processor from 4 to 5.
71050725 2015 206 11:15 2 9:01:01 Timing EC-1073 (changing 1PPS source) was installed into the system. The configuration flags were incremented (on day 208 00:20) from 2 to 3 on the controller and 3 to 4 on the processor.
71050725 2015 203 08:15 0 99 System back up, the high voltage cable connector was replaced to clear the problem with the high voltage power supply voltage and current fluctuations.
71050725 2015 202 18:35 0 99 System down due to High voltage power supply voltage and current fluctuation.
71050725 2015 200 00:25 0 99 Original LRC reinstalled back into the system: incremented the configuration flags on the controller from 1 to 2 and the processor from 2 to 3 for MOBLAS 7.
71050725 2015 194 22:00 2 9:01:01 Added a cable delay to the XLDC +7230nsec to bring the GPS-RUB to -7.061us and the STA-RUB to 7.078us.
71050725 2015 194 21:12 2 99 Installed HP5370B counter G78681: configuration flags were incremented on the controller from 0 to 1 and the processor from 1 to 0 for the test of.
71050725 2015 181 11:11 0 99 Network spare LRC chassis installed into the system for testing: incremented the configuration flags (day 182 02:47) on the processor computer from 0 to 1 and controller from 9 to 0 for the spare LRC installation.
71050725 2015 172 23:58 2 6.01.01 Replaced the MCP due to weak returns; incremented the configuration flag (day 173 02:31) on the processor from 9 to 0 for the replacement of the MCP and receive cable.
71050725 2015 170 09:30 0 99 High voltage power supply voltage and current is fluctuating, replaced bad high voltage cable connector.
71050725 2015 156 11:12 2 9:01:01 Engineering working on timing rack, replaced CNS, G77945 to G77979; incremented the configuration (at 20:19) on the processor computer from 8 to 9 for the replacement of the CNS unit.
71050725 2015 153 10:55 0 99 Replaced Ethernet switch under floor with new gigabit Ethernet switch model G5108 NTGEAR, downlink 94.79mbbps, uplink 97.72mbps.
71050725 2015 137 09:15 0 99 Using timing to verify the spare LRC chassis, had to reset timing TCG and CNS to XLDC reading changed by 30nsec.
71050725 2015 118 15:45 2 99 Removed MET4 000829 and installed MET4 2378313.
71050725 2015 049 20:15 2 6.01.01 Replaced the MCP gate cable, the one in the system failed.
71050725 2014 300 02:53 2 6.01.01 Replaced the MCP; the one in the system failed; config flag on process incremented from 6 to 7.
71050725 2014 254 02:11 2 99 Removed M04297 counter from M7 after testing and reinstalled M7 counter G78769; config flag on process incremented from 5 to 6.
71050725 2014 224 21:17 0 99 Timing jumped, Sta-Rub was 7.160us now 7.192us; GPS-Rub was -7.202us now -7.103us; reason unknown.
71050725 2014 210 18:00 2 99 Installed M04297 counter into M7 for testing; config flag on process incremented from 4 to 5.
71050725 2014 209 13:00 2 99 Removed G71282 counter and reinstalled M7 counter; config flag on process incremented from 3 to 4.
71050725 2014 202 13:00 2 99 LRC chassis false stop adjusted from one microseconds to two microseconds.
71050725 2014 200 13:00 2 9 IRIG B sync out of TCG problems.
71050725 2014 189 13:00 0 5 Laser adjusted and bore-sighted on the ground target.
71050725 2014 184 13:00 2 99 Installed G71282 counter; config flags reset from 2 to 3 on processor computers.
71050725 2014 182 13:00 0 99 Replaced two bad cables for door interlock.
71050725 2014 167 13:00 2 11 Changed out transceiver on Radar to eliminate tuning errors.
71050725 2014 166 02:45 2 99 Installed repaired G78769; config flags reset 1 to 2.
71050725 2014 166 01:34 2 99 Installed SCNS0003190; removed G78769; config flags reset 0 to 1.
71050725 2014 163 13:00 2 99 Installed G77341; removed G77986 bad power supply.
71050725 2014 161 13:00 2 99 Installed G71282; returned G78769; config flags reset 9 to 0.
71050725 2014 160 13:00 2 99 Removed G77986; installed G77341.
71050725 2014 152 13:00 0 99 Installed repaired OAM daylight filter switch.
71050725 2014 150 13:00 2 99 Repaired spare LRC and ARSU Chassis.
71050725 2014 147 13:00 0 99 Laser Interlock Door Failure; installed spare cards.
71050725 2014 139 16:00 2 6.01.01 Installed spare MCP into M7 system due to gating issues; config flags reset from 7 to 8.
71050725 2014 137 16:00 2 6.01.01 Replaced PMT; using original cable.
71050725 2014 135 13:00 2 6 Replaced OAM and Gating cables.
71050725 2014 125 13:00 2 99 Spare counter M04263 installed for Network Testing and config flag set to 7 from 6 (G78769 removed).
71050725 2014 122 10:00 2 6 Receive cable bad; replaced.
71050725 2014 081 00:00 1 9.01.01 XL-DC replaced and removed XL-DC battery backup.
71050725 2014 065 22:00 2 6.01.01 installed new Photek MCP.
71050725 2013 305 00:00 1 9.01.01 Replaced A10 in the TCG.
71050725 2013 143 17:00 1 9.01.01 XLDC 000271 Removed 000149 Installed.
71050725 2013 115 04:06 2 99 Installed 5370B counter G78769
71050725 2013 071 12:58 2 6 Receive Cable changed
71050725 2013 022 08:05 2 99 Installed 5370B counter G78768
71050725 2012 354 00:00 1 9.01.01 New Truetime installed (G77318).
71050725 2012 348 22:10 1 9.01.01 Replaced Truetime XL-DC Freq. Standard.
71050725 2012 347 10:10 2 09.01 XL - DC G000271 was installed
71050725 2012 296 00:00 1 9 A10 card replaced in TCG to fix strobe fault.
71050725 2012 263 02:10 2 6.01.01 Installed MCP daylight filter assy.
71050725 2012 263 02:05 2 9 Installed new TCG G77986.
71050725 2012 261 17:00 1 12.02.01 MET4 117635 Removed MET4 115135 Installed.
71050725 2012 101 05:28 2 99 Counter return to orginal configuration
71050725 2012 096 04:12 2 99 Counter G78409 installed
71050725 2012 089 03:57 2 99 Counter G71203 installed
71050725 2012 082 05:15 2 99 Counter G71204 installed
71050725 2012 069 04:30 2 99 Counter G77376 installed
71050725 2012 068 06:20 2 99 Counter G78762 installed
71050725 2012 044 04:16 2 99 Low loss cable changed
71050725 2011 332 13:99 2 99 Reversed cables for M7 SLR tracking.
71050725 2011 332 11:00 2 9 Removed W518 start cable from 5370B transmit delayed counter. Installed start diode to cable.
71050725 2011 237 13:00 1 6.01.01 Removed RUB 100s and installed Phillips Scientific #460.
71050725 2011 235 14:51 2 99 M5.8 Earthquake about 94 miles from Greenbelt near Mineral, VA - survey post earthquake shows no conclusive change
71050725 2011 153 02:10 2 99 Counter G78769 installed
71050725 2011 143 08:29 2 99 Took card from G77977 and installed into G78768.
71050725 2011 141 09:20 2 99 Counter G71279 installed
71050725 2011 140 10:18 2 99 Counter G78768 installed
71050725 2011 136 12:17 2 99 Counter G78769 installed
71050725 2011 130 08:55 2 99 Counter G78793 installed
71050725 2011 124 11:35 2 99 Counter G71283 installed
71050725 2011 119 01:41 2 99 Counter G78777 installed
71050725 2011 115 06:12 2 09.01 Changed XL-DC antenna
71050725 2011 111 09:22 2 99 Installed Counter
71050725 2011 109 01:35 1 9 Replaced G77390 with CNS Clock G77945.
71050725 2011 108 00:50 1 9 CNS Clock G77945 replaced with G77390.
71050725 2011 097 15:00 1 12.02.01 MET4 G71224 Installed T = 13.2; H = 54.3 and B = 1011.17
71050725 2011 097 14:27 1 12.02.01 MET4 G78677 replaced T = -0.12.4 H = 56.4; B = 1010.67
71050725 2011 081 00:00 2 99 HP-5370B Counter G78023 Removed and G78775 Installed
71050725 2011 066 11:06 2 99 Installed G78769 counter into system for satellite tracking.
71050725 2011 062 06:41 2 99 Installed counter G78023 into system.
71050725 2011 062 05:01 2 99 Installed counter G78770 into system.
71050725 2011 062 03:36 2 99 Installed counter G71283 into system.
71050725 2011 059 04:01 2 99 Installed counter G78793 into system.
71050725 2011 053 06:58 2 99 Installed counter G78775 into system.
71050725 2011 047 00:01 2 99 Returned M7 counter G78769 back into system.
71050725 2011 020 08:00 2 99 Installed G78761.
71050725 2011 020 07:03 2 99 Installed G71203.
71050725 2011 020 05:32 2 99 Installed G78777.
71050725 2010 344 13:00 1 5.01.01 10PPS problem found to be replaced oscillator last month had less output power. Removed some attenuation from photo diode (went from .5 @ 5PPS and .38 @ 10pps to 2.3 @ 5PPS and 1.4 @ 10PPS).
71050725 2010 315 21:00 1 6.01.01 Problem with PMT discovered elbow HV connector pin out of socket re-assembled all okay.
71050725 2010 278 05:00 2 99 G78557 counter not working; install s/n: 2410A00829.
71050725 2010 265 01:00 1 6.01.01 Installed M6 MCP; it is operational.
71050725 2010 224 15:30 1 6.01.01 Installed ITT MCP; gating module; and delay box for testing.
71050725 2010 223 13:30 1 6.01.01 Installed new PMT.
71050725 2010 197 05:00 1 6.01.01 Installed M4 MCP.
71050725 2010 149 02:17 2 99 Counter changes(G78501; and ECN 095894).
71050725 2010 147 00:00 2 99 Multiple counter changes throughout the day (G78501; G78380; G78024; G78557; G78501; then G78557)
71050725 2010 143 11:53 2 99 Installed G78501 counter into the system.
71050725 2010 143 06:54 2 99 Installed G78612 counter into system.
71050725 2010 138 05:30 2 99 Installed M7 spare counter (G78409) into system for two passes; then returned M7 original back into system (G78557)
71050725 2010 132 10:36 1 9 10MHz is removed from back of 5370B counters for future testing of spare 5370Bs.
71050725 2009 334 00:00 1 9.01.01 Failed XL-DC unit (G77108) replaced with G77401.
71050725 2009 302 00:00 1 9.01.01 Changed out GPS antenna.
71050725 2009 297 00:00 1 9 Replaced CNS clock.
71050725 2009 274 14:36 1 12.02.01 MET3 Removed MET4 74875 Installed
71050725 2009 254 12:45 2 99 Installed backup counter to compare with operational unit to troubleshoot 10PPS range word error.
71050725 2009 252 14:00 2 99 Had to modify the boards (switched a wire on all 4 boards).
71050725 2009 192 00:00 2 99 Switched out LRC board for 10PPS.
71050725 2009 136 00:00 0 99 Troubleshot discriminator/NIMBIN issues (removing/installing cards/power supply).
71050725 2009 135 00:00 0 99 Lost -6 volts; replaced receive discriminator.
71050725 2009 134 00:00 0 99 NIMBIN Power Supply replaced.
71050725 2009 113 00:00 0 99 Replaced NIMBIN with one on slides.
71050725 2009 093 00:00 1 99 Installed M4 discriminator.
71050725 2009 089 00:00 1 99 Changed operational discriminator with spare from GCP.
71050725 2009 078 00:00 1 6 Removed old receive cable and ran the spare receive cable; installed new processor computer
71050725 2009 072 13:00 1 5.0 Laser EC-010-09-2007 (chg-1) Laser Chiller Upgrade
71050725 2009 065 00:00 1 6.01.01 Replaced MCP Cable to correct system delay.
71050725 2009 041 00:00 2 9 Stability tests showed system delay drifts; replaced 5 MHz cable from Dist. Amp to 5370B.
71050725 2009 037 13:00 2 99 Computer EC-016-04-2009 (chg-1) Processor Computer Upgrade
71050725 2008 293 13:00 1 5.01 Laser EC-015-10-2008 (rev 1) Laser Dye Cell Replacement (7/20/09)
71050725 2008 195 00:00 0 99 Changed circuit breaker and moved from phase "C" to phase "A".
71050725 2008 151 13:00 1 99 Removed Td connections at patch for signals going to 53131A counter for extended testing of GPS v Rb.
71050725 2008 143 13:00 1 99 Td connections at patch for signals going to 53131A counter for extended testing of GPS v Rb.
71050725 2008 115 00:00 1 9 Installed cable on 5370B 10MHz to drive 53132A counter being used for LRO M7 timing study.
71050725 2008 079 00:00 2 9 Switched 5370B; new unit G78557 (old unit G78382).
71050725 2008 073 00:00 2 12.01.02 Replaced MET3 with spare from GCP (G78390).
71050725 2008 064 00:00 2 12.01.01 G78440 MET-3 replaced with G78502 MET-3.
71050725 2007 273 13:00 2 99 Computer EC-011-11-2007 (chg-1) Controller Computer Upgrade
71050725 2007 258 13:00 0 99 UPS EC-009-09-2007 APC UPS Upgrade Installation
71050725 2007 226 00:00 2 9 5370B reading error; removed and checked system configured internally; re-installed after checking okay offline; worked fine after re-installation).
71050725 2007 169 03:00 2 9 Installed spare 5370B (G78382).
71050725 2007 043 00:00 2 99 G77973 model 53131A re-installed.
71050725 2007 008 00:00 2 99 G77342 model 5334B counter replaced with G77618 model 53131A. Replaced G77392 model 53131A with G77973 model 53131A. M7 counters in calibration.
71050725 2006 150 07:20 2 9 Problem with Radar was traced to TCG; bad cap A10 C1. Repaired.
71050725 2006 121 00:36 1 12.03.01 Changed out the MET3 unit.
71050725 2006 006 00:00 1 12.03.01 New calibrated MET-3 installed & verified.
71050725 2005 131 00:00 2 99 5370B removed to connect T-4s 5370B. Cable not connected back up when done.
71050725 2004 328 10:04 2 99 Removed 5370B L-3111 (G77264) replaced with G78380.
71050725 2004 279 00:00 1 9 Replaced C1 of U1:A board 10 of TCG.
71050725 2004 265 23:57 1 9 Found BNC connector on TCG bad; repaired.
71050725 2003 343 21:00 1 99 L-5848 removed.
71050725 2003 318 13:00 0 99 Connected a 1PPS cable to the TCG for the GUTS system.
71050725 2003 272 13:00 1 12.01.01 New Met sensor installed; hung-up OAM; added sync cable for the GUTS system to M7 controller.
71050725 2003 210 19:00 1 9 Returned the TIU counter.
71050725 2003 197 21:00 1 6.01.01 Troubleshot the MCP tube; found it bad/removed.
71050725 2003 195 13:00 1 6 Changed receive cable.
71050725 2003 186 13:00 1 6.01.01 New PMT (21021004) and new CFD (867) installed.
71050725 2003 161 21:30 1 6.01.01 M7 MCP & CFD re-installed (CFD #867/PMT #21021004 replaced with CFD #265 and PMT #43020531).
71050725 2003 121 20:09 1 6.01.01 New MCP/CFD combination installed.
71050725 2003 093 05:00 2 99 Removed L-5848 and replaced with L-4970 after original stopped counting; reinstalled original after GCP checked it out and founding nothing wrong.
71050725 2003 084 17:51 1 6.01.01 Installed M6 Photek PMT and CFD for testing.
71050725 2003 056 02:00 1 6 Change new receive cable to PP1.
71050725 2003 044 15:00 1 9 TIU start/stop/BNC Adapter replaced.
71050725 2003 032 21:00 1 99 Installed L-3398 CFD.
71050725 2003 016 19:16 1 99 Installed L-5848.
71050725 2003 015 13:00 2 99 Multiple swaps in/out of 5370B; original unit re-installed.
71050725 2003 012 13:00 2 99 Replaced 5370B L-3111.
71050725 2002 329 20:43 1 6.01.01 M4 MCP/CFD equipment removed and M7 version installed.
71050725 2002 326 13:00 1 6.01.01 New photek PMT installed s/n: 42020802 for system testing. M4 CFD was left in the system.
71050725 2002 304 14:11 1 6.01.01 Installed Photek MCP/PMT s/n: 42020531 to conduct testing of the tube and CVD for M4 (M7s CFD L-5167 removed to install M4s unit).
71050725 2002 236 13:00 1 6 New receive cable broken replaced with old cable.
71050725 2002 236 13:00 1 6 New receive cable put back in.
71050725 2002 225 13:00 1 6 all Installed s/n 43020531 MCP.
71050725 2002 224 13:00 1 6.01.01 Switched back to ITT MCP.
71050725 2002 220 18:18 1 6.01.01 Installed s/n 23020625 Photek MCP.
71050725 2002 217 22:57 1 6.01.01 Photek PMT swapped for new unit (s/n 43020531).
71050725 2002 214 00:00 1 6.01.01 New Photek PMT installed
71050725 2002 202 23:15 1 99 Swapped out discriminators L-3398 with L-5167 (L-3398 was noisy).
71050725 2002 189 13:00 1 99 NIMBIN power supply replaced.
71050725 2002 171 20:53 1 9.01.01 LX-DC removed and replaced with L-5803
71050725 2002 119 17:34 1 9 HP5087A replaced (L-4011 swapped with L-4088).
71050725 2002 028 13:00 2 6.01.01 MCP from GCP testing.
71050725 2002 014 00:00 1 12.01.01 Replaced MET 3 sensor with newly NIST certified unit.
71050725 2001 352 13:00 1 6.01.01 ETN 026-628 Burle 85104 MCP-PMT Testing installed.
71050725 2001 312 13:00 2 6.01.01 New PMT cable installed.
71050725 2001 310 14:38 2 9.01.01 Pulled the XL-DC L-5536 (unit to be refurbished/replaced) and replaced with s/n: 010514244.
71050725 2001 307 04:04 0 9.01.01 Batteries replaced in timing standby power supply (L-4296). Power cord replaced for the Truetime GPS & Freq. Standard.
71050725 2001 284 21:26 1 9.01.01 Standby Power Supply issues repaired. Attempt to replace XL-DC but spare was unable to lock so original was returned to system.
71050725 2001 284 13:00 1 9.01.01 Spare frequency receiver installed.
71050725 2001 267 13:00 2 6.01.01 Installed MCP from GCP (Hollis).
71050725 2001 255 13:00 2 6.01.01 PMT swapped out.
71050725 2001 245 17:00 2 6.01.01 MCP down; repaired; system back operational.
71050725 2001 243 21:00 2 6.01.01 MCP problem; changed out signal cable.
71050725 2001 236 05:00 2 99.01.01 Installed new Tac324 program.
71050725 2001 035 21:00 1 6.01.01 MCP removed/eyepiece is in.
71050725 1999 309 16:00 1 9 Replaced the TCG Power Supply.
71050725 1999 228 13:00 1 11 Radar Installed Radar Safety Upgrade 20 PPS @50% duty cycle mode to A10 card of TCG.
71050725 1999 084 05:00 2 6 repaired MCP cable and re-tied cables to reduce strain.
71050725 1999 063 16:00 1 9.01.01 GPS Trimble GPS (L-4643) and its antenna removed.
71050724 1999 049 17:00 2 6 Replaced receive PMT cable.
71050724 1999 035 23:00 2 9 New cables installed to OAM & PMT.
71050724 1998 339 13:00 1 9 Connected 10MHz line from distribution amplifier to TCG.
71050724 1998 315 17:00 1 9.01.01 GPS and Frequency standard switch installed.
71050724 1998 194 13:00 1 9.01.01 Swapped Cesium L-4256 with L04296.
71050724 1997 315 13:00 1 12.03.01 Barometer now working; replaced multiplexer board.
71050724 1997 163 05:00 1 99 Pulse Generator card removed from NIMBIN.
71050724 1997 161 13:00 1 9.01.01 GPS equipment removed.
71050724 1997 150 13:00 2 6.01.01 MCP Receive cable replaced (broken).
71050724 1997 093 17:00 0 99 Interchanged U7 & J9 in ARSU to eliminate relay chattering.
71050724 1997 086 17:00 1 5 Krytron change out
71050724 1997 084 13:00 1 99 ARSU installed similar to Moblas-6 unit.
71050724 1996 318 21:00 2 5.01 Completed installation of MCP cable with new range counter.
71050724 1996 250 21:00 2 9.01.01 Installed MLRO Timing cables to/from TCG and distribution amplifier.
71050724 1996 117 14:00 1 5.01 Repaired PMT HV Supply Monitor Circuit.
71050724 1996 096 16:00 0 99 Replaced defective switch on ARSU.
71050724 1996 089 16:00 2 5.01 MCP & HSLR discriminator returned; re-cabled integrator gating to match block diagram.
71050724 1996 088 16:00 0 9.01.01 Replaced 20pps cable from Time Code Generator to Encoders cable #W315
71050724 1996 072 06:00 0 99 Corrected wiring error in LRC. Correct window now triggering the A-Scope
71050724 1996 023 05:00 0 9.01.01 Cesium 161B replaced with cesium 870. New STA-Cs = 198.5 microsecs. TCG Leap year switch activated.

Find more topics on the central web site of the Technical University of Munich: www.tum.de